OpenStudy (anonymous):

Sove the IVP xy' + y = 0 y(4) = 6

OpenStudy (anonymous):

@zepdrix I did this: \[y'=-y/x\\y'=-u\\ y=-u^2/2 + c\] Did I make a mistake with the integration part?

zepdrix (zepdrix):

Oh did they want you to make a substitution again?

zepdrix (zepdrix):

Yah looks ok so far. If the instructions didn't ask you to make a substitution, I would avoid doing so. Because you're going to end up with an answer of \(\large y=f(x,y)\) x's and y's on the right. And this is a nice easy problem that can be solved in terms of \(\large y=f(x)\)

sam (.sam.):

or just use reverse product \[xy' + y = 0\] \[\frac{d}{dx}(xy)=0\] \[xy=\int\limits 0dx\]

zepdrix (zepdrix):

\[\large y'=-\frac{y}{x}\qquad\rightarrow\qquad \frac{1}{y}dy=-\frac{1}{x}dx\]

zepdrix (zepdrix):

Ah yes, that works nicely too c: what sam did.

OpenStudy (anonymous):

Oh I see, I should rather just do the normal separable method thing...

OpenStudy (anonymous):

so then I would get\[\ln|y| = -\ln|x| + c_1\\y=-x+c_2\] And with the initial values c_2 = 10 and y=-1+10 ?

OpenStudy (anonymous):

ah y=-x+10

zepdrix (zepdrix):

Woops after you integrated, let's be careful with the next step.

zepdrix (zepdrix):

\(\large \ln y=-\ln x+c \qquad \rightarrow \qquad \ln y=\ln\left(\dfrac{1}{x}\right)+c\) Exponentiating both sides gives us,\[\large y=e^{\left(\ln\left(\dfrac{1}{x}\right)+c\right)} \qquad \rightarrow \qquad y=C\frac{1}{x}\]

zepdrix (zepdrix):

That unknown constant is now a factor, it's no longer being added, due to the exponentiation. I can show another step in there if you're confused by the last part.

OpenStudy (anonymous):

Thanks, I'm not confused yet haha. I will be on the lookout for this kind of ln thing, I tend to do what I just did...

zepdrix (zepdrix):

Yah it's a very easy mistake to make :) Make sure you exponentiate EVERYTHING on the right! heh

OpenStudy (anonymous):

Thanks! In this problem, would we get y=24/x?

zepdrix (zepdrix):

Ah yes, that sounds right.

OpenStudy (anonymous):

@zepdrix I did another problem which led me to this: \[\ln|y|=2x^2-\ln|\cos x| +C\] After this is exponentiated, how do I simplify the equation?

zepdrix (zepdrix):

\[\large y=e^{\left(2x^2-\ln|\cos x| +c\right)}\] \[\large y=e^{(2x^2-\ln(\cos x))}\cdot e^c\] \[\large y=Ce^{(2x^2-\ln(\cos x))}\] Hmm like that I suppose.

OpenStudy (anonymous):

I think I know... I should try to combine the ln's? So,\[\ln|y| + \ln|\cos x|=2x^2 +C\\ \ln|y \cos x|=2x^2 + C \\ y \cos x = c_2 e^{2x^2} \\ y = c_2\frac{e^{2x^2}}{\cos{x}}\]?

zepdrix (zepdrix):

Oh Yes i missed a step in there, my bad. \[\large y=Ce^{(2x^2+\ln(\sec x))}\]\[\large y=Ce^{2x^2}\cdot e^{\ln \sec x}\]\[\large y=C (\sec x) e^{2x^2}\] Your method seems to have worked also :) Nice

OpenStudy (anonymous):

Thanks zep!